# intersection.bib

### v2

@@ 3 Bounds for the furthest distance a car can come from
For sufficiently large $n,\ell$ there exist constants $C,c>0$ (depending only on $\epsilon$) such that
@@
This lemma is already proved in [BBB23, Lemma 3.1], the Reader can refer to this proof. We briefly describe the idea here. We consider two points $a$ and $b$ on $\mathcal{L}_{-n-\lceil\frac{n}{100k_{i,j}}\rceil}$ which are $\lceil 2\ell n^{1/96}n^{2/3}\rceil$ distance away (in the space direction) from the end points of $R_{i}^{\ell}$. Now, by transversal fluctuation estimate we can say that $\Gamma_{a}^{\beta_{i}}$ and $\Gamma_{b}^{\alpha_{i}}$ stays always $\lceil\ell^{1/96}n^{2/3}\rceil$ distance away (in the space direction) from the rectangle $R_{i}^{\ell}$ with probability at least $1-e^{-c\ell^{3/96}}$. On this event, every geodesics starting in $R_{i}^{\ell}$ will be sandwiched between $\Gamma_{a}^{\beta_{i}}$ and $\Gamma_{b}^{\alpha_{i}}$. So, on this event by planarity argument the event $\mathcal{G}_{i}$ will imply large transversal fluctuation of either $\Gamma_{a}^{\beta_{i}}$ or $\Gamma_{b}^{\alpha_{i}}$. This gives us the desired upper bound. ∎
@@
Note that on the second event in the right hand side there are more than $\ell^{1/3}$ distinct equivalence classes of geodesics starting from $\mathcal{J}_{i}^{\prime}$ and ending at $\mathcal{J}_{i}^{\prime\prime}$. Using Proposition 3.4, for $\ell<n^{0.03}$ sufficiently large (depending on $\epsilon$) we have
@@
Same argument shows for $\ell<n^{0.03}$ sufficiently large (depending on $\epsilon$) we have
@@
$\mathbb{P}\left(\max I_{u,v}\geq\frac{\ell^{1/3}n\log|k_{i,j}|}{k_{i,j}^{3}}\right)\leq\mathbb{P}(\mathcal{F}_{i}\cup\mathcal{G}_{i}\cup\mathcal{F}_{j}\cup\mathcal{G}_{j})+$
($\mathbb{P}\left(\bigl{\{}\max I_{u,v}\geq\frac{\ell^{1/3}n\log|k_{i,j}|}{k_{i,j}^{3}}\bigr{\}}\cap(\mathcal{F}_{i}\cup\mathcal{G}_{i}\cup\mathcal{F}_{j}\cup\mathcal{G}_{j})^{c}\right).$
@@
On the event $\bigl{\{}\max I_{u,v}\geq\frac{\ell^{1/3}n\log|k_{i,j}|}{k_{i,j}^{3}}\bigr{\}}\cap(\mathcal{F}_{i}\cup\mathcal{G}_{i}\cup\mathcal{F}_{j}\cup\mathcal{G}_{j})^{c}$, we want to estimate the maximum size of the intersection of geodesics inside $V$ that are starting from $\mathcal{J}_{i}^{\prime}$ (resp. $\mathcal{J}_{j}^{\prime}$) and ending at $\mathcal{J}_{i}^{\prime\prime}$ (resp. $\mathcal{J}_{j}^{\prime\prime}$). For this we have the following lemma.
Lemma 3.14.
[BBB23, Lemma 3.2, Lemma 3.3] Consider any geodesic $\Gamma_{i}$ (resp. $\Gamma_{j}$) starting from $\mathcal{J}_{i}^{\prime}$ (resp. $\mathcal{J}_{j}^{\prime}$) and ending at $\mathcal{J}_{i}^{\prime\prime}$ (resp. $\mathcal{J}_{j}^{\prime\prime})$. If $I_{\Gamma_{i},\Gamma_{j}}$ denote the intersection size of $\Gamma_{i}$ and $\Gamma_{j}$ inside $W$ then for $\ell$ and $n$ large enough there exist $C,c>0$ such that we have
Hence, we have that there exist constants $C,c$ (depending on $\epsilon$) such that for large enough $\ell$ and $n$
($\mathbb{P}\left(\bigl{\{}\max I_{u,v}\geq\frac{\ell^{1/3}n\log|k_{i,j}|}{k_{i,j}^{3}}\bigr{\}}\cap(\mathcal{F}_{i}\cup\mathcal{G}_{i}\cup\mathcal{F}_{j}\cup\mathcal{G}_{j})^{c}\right)\leq Ce^{-c\ell^{1/3}}.$
Hence, combining the above we get that there exists a constant $C>0$ (depending on $\epsilon$) such that for all large enough $n$ we have
@@
hence, from (3.10) we have there exists $C_{2}>0$ such that for sufficiently large $n$
@@
We now go back to (3.7). From (3.9) and (3.15) we have there exists constants $c_{1},C_{2}>0$, large $M$ such that for sufficiently large $n$(all depending only on $\epsilon$)
@@
As $C,c$ depend only on $\epsilon$ (as we have chosen all the interval length to be $\frac{1}{n^{1/3}}$) we can choose $M$ large enough so that there exists $c>0$ such that for sufficiently large $n$

### v3

@@ 3 Bounds for the furthest distance a car can come from
Figure 6: The figure shows how to obtain an upper bound for $\mathbb{P}(\mathcal{F}_{i})$ when $\theta_{i}>\theta_{j}>\frac{\pi}{4}.$ Although $W_{i,j}$ is aligned in the direction $\theta_{j}$, if for some $u\notin\overline{\mathsf{R}_{i}^{\ell}}$ and $\gamma_{i}\in A_{i}$ intersect $W_{i,j}$ then it will have large transversal fluctuation of order at least $\sim(\ell^{1/96}-1)n^{2/3}$ as shown in the figure. Thus, $\mathbb{P}(\mathcal{F}_{i})$ will be exponentially small in $\ell$. The other cases of $\theta_{i},\theta_{j}$ follow similarly.
Proof.
We assume the case when $\theta_{i}>\theta_{j}>\frac{\pi}{4}$. The other cases will follow similarly. The upper bound for $\mathbb{P}(\mathcal{F}_{j})$ is immediate from Proposition 6.3 as the parallelogram $W_{i,j}$ is aligned in the direction $\theta_{j}.$ For the event $\mathcal{F}_{i},$ first define $\overline{\mathsf{R}_{i}^{\ell}}$ as the parallelogram whose opposite sides lie on $\mathcal{L}_{-n-\lceil\frac{n}{100k_{i,j}}\rceil}$ (resp. $\mathcal{L}_{\lceil\frac{n}{100k_{i,j}}\rceil}$), each of length $\ell^{1/96}n^{2/3}$ with midpoints at the intersections of $y=\tan(\theta_{i})x$ and $\mathcal{L}_{-n-\lceil\frac{n}{100k_{i,j}}\rceil}$ (resp. $\mathcal{L}_{\lceil\frac{n}{100k_{i,j}}\rceil}$, see Figure 6). Although $W_{i,j}$ is aligned in the direction $\theta_{j}$ which is different from $\theta_{i}$, we now show that it is still properly contained in the parallelogram $\overline{\mathsf{R}_{i}^{\ell}}$ for sufficiently large $\ell$.
To see this, it is enough to verify the containment on the lines $\mathcal{L}_{\lceil\frac{n}{100k_{i,j}}\rceil}$ and $\mathcal{L}_{-\lceil\frac{n}{100k_{i,j}}\rceil}$. As these lines have direction $\frac{3}{4}\pi$, containment of $W_{i,j}$ in $\overline{\mathsf{R}_{i}^{\ell}}$ is implied by the corresponding ordering of the horizontal coordinates for the intersections of the respective parallelograms with these two lines. We show exactly this below. By definition, the sides of $W_{i,j}$ on these lines have midpoints that come from the system
$y&=\tan(\theta_{j})x\\ x+y&=\lceil\frac{n}{100k_{i,j}}\rceil\end{aligned}\right\}\qquad\text{and}\qquad\left.\begin{aligned} y&=\tan(\theta_{j})x\\ x+y&=-\lceil\frac{n}{100k_{i,j}}\rceil\end{aligned}\right\}.$
That is, the upper side of $W_{i,j}$, of side length $n^{2/3}$, spans through horizontal coordinates
$x\in\Bigl{(}\frac{\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{j}}-\frac{1}{2\sqrt{2}}n^{2/3},\,\frac{\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{j}}+\frac{1}{2\sqrt{2}}n^{2/3}\Bigr{)}$
on $\mathcal{L}_{\lceil\frac{n}{100k_{i,j}}\rceil}$, while the lower side spans
$x\in\Bigl{(}\frac{-\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{j}}-\frac{1}{2\sqrt{2}}n^{2/3},\,\frac{-\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{j}}+\frac{1}{2\sqrt{2}}n^{2/3}\Bigr{)}$
on $\mathcal{L}_{-\lceil\frac{n}{100k_{i,j}}\rceil}$. To simplify notation we increase this for all large $n$ by dropping the constant factors on $n^{2/3}$.
The parallelogram $\overline{\mathsf{R}_{i}^{\ell}}$ has side length at least $2\sqrt{2}\ell^{1/97}n^{2/3}$ for large $\ell$ (i.e., by slightly decreasing the power on $\ell$ we don’t need to carry constant factors below). We then conclude, similarly to how we did for $W_{i,j}$, that the intersection of $\overline{\mathsf{R}_{i}^{\ell}}$ with the line $\mathcal{L}_{\lceil\frac{n}{100k_{i,j}}\rceil}$ contains the horizontal coordinates
$x\in\Bigl{(}\frac{\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{i}}-\ell^{1/97}n^{2/3},\,\frac{\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{i}}+\ell^{1/97}n^{2/3}\Bigr{)}$
and, with the line $\mathcal{L}_{-\lceil\frac{n}{100k_{i,j}}\rceil}$, the horizontal coordinates
$x\in\Bigl{(}\frac{-\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{i}}-\ell^{1/97}n^{2/3},\,\frac{-\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{i}}+\ell^{1/97}n^{2/3}\Bigr{)}.$
In view of the above, to show containment it is enough to demonstrate that
$\Bigl{(}\frac{\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{j}}-n^{2/3},\,\frac{\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{j}}+n^{2/3}\Bigr{)}\subseteq\Bigl{(}\frac{\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{i}}-\ell^{1/97}n^{2/3},\,\frac{\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{i}}+\ell^{1/97}n^{2/3}\Bigr{)}$
and
$\Bigl{(}\frac{-\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{j}}-n^{2/3},\,\frac{-\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{j}}+n^{2/3}\Bigr{)}\subseteq\Bigl{(}\frac{-\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{i}}-\ell^{1/97}n^{2/3},\,\frac{-\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{i}}+\ell^{1/97}n^{2/3}\Bigr{)}.$
Due to symmetry, these four inequalities reduce to the below two:
$\frac{\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{j}}-n^{2/3}$ $\geq\frac{\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{i}}-\ell^{1/97}n^{2/3},$ (3.10)
$\frac{\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{j}}+n^{2/3}$ $\leq\frac{\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan\theta_{i}}+\ell^{1/97}n^{2/3}.$ (3.11)
(3.10) is easy in the case $\theta_{i}>\theta_{j}>\frac{\pi}{4}$ since for large $\ell$ the terms on the two sides of the inequality dominate each other pairwise.
(3.11) is more involved. Rearranging, we need to show that
$\lceil\frac{n}{100k_{i,j}}\rceil\Bigl{(}\frac{1}{1+\tan\theta_{j}}-\frac{1}{1+\tan\theta_{i}}\Bigr{)}$ $\leq\bigl{(}\ell^{1/97}-1\bigr{)}n^{2/3},$
$\lceil\frac{n}{100k_{i,j}}\rceil\cdot\frac{\tan\theta_{i}-\tan\theta_{j}}{(1+\tan\theta_{j})(1+\tan\theta_{i})}$ $\leq\bigl{(}\ell^{1/97}-1\bigr{)}n^{2/3}.$
Dropping the second denominator can only increase the left-hand side. Recalling $1\leq j<i\leq\bigl{[}\frac{n^{1/3}}{2M}\bigr{]}$ and $k_{i,j}=M(i-j)$, we also see that the first fraction is at least in the order of $n^{2/3}$. Hence we can also get rid of the integer part by adding in a factor of 2. It is therefore enough to show
$\frac{n}{50k_{i,j}}\cdot(\tan\theta_{i}-\tan\theta_{j})\leq\bigl{(}\ell^{1/97}-1\bigr{)}n^{2/3}.$
Recall that the angles are each in the range $(\frac{\pi}{4},\,\frac{\pi}{2}-\varepsilon)$. The derivative of $\tan$ is bounded here, uniformly in all parameters except $\varepsilon$. Thus, with $k_{i,j}=M(i-j)$ and the definition $\theta_{i}=\frac{Mi(\frac{\pi}{2}-2\varepsilon)}{n^{1/3}}$, we now need to prove
$\text{const}\cdot\frac{n}{i-j}\cdot\frac{i-j}{n^{1/3}}\leq\bigl{(}\ell^{1/97}-1\bigr{)}n^{2/3},$
which is trivially true for large $\ell$.
Now we show that if there is a geodesic starting outside of $R_{i}^{\ell}$ and taking angle in $A_{i}$ intersects $W_{i,j}$ then it will imply large transversal fluctuation. Note that from the above discussion, the end points of $W_{i,j}$ are (see Figure 6)
$a:$ $=\left(\frac{-\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan(\theta_{j})}-\frac{1}{2\sqrt{2}}n^{2/3},\frac{-\lceil\frac{n}{100k_{i,j}}\rceil\tan(\theta_{j})}{1+\tan(\theta_{j})}+\frac{1}{2\sqrt{2}}n^{2/3}\right),$
$b:$ $=\left(\frac{-\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan(\theta_{j})}+\frac{1}{2\sqrt{2}}n^{2/3},\frac{-\lceil\frac{n}{100k_{i,j}}\rceil\tan(\theta_{j})}{1+\tan(\theta_{j})}-\frac{1}{2\sqrt{2}}n^{2/3}\right),$
$c:$ $=\left(\frac{\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan(\theta_{j})}-\frac{1}{2\sqrt{2}}n^{2/3},\frac{\lceil\frac{n}{100k_{i,j}}\rceil\tan(\theta_{j})}{1+\tan(\theta_{j})}+\frac{1}{2\sqrt{2}}n^{2/3}\right),$
$d:$ $=\left(\frac{\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan(\theta_{j})}+\frac{1}{2\sqrt{2}}n^{2/3},\frac{\lceil\frac{n}{100k_{i,j}}\rceil\tan(\theta_{j})}{1+\tan(\theta_{j})}-\frac{1}{2\sqrt{2}}n^{2/3}\right).$
Similarly, the corresponding vertices of $R_{i}^{\ell}$ restricted between $\mathcal{L}_{-\lceil\frac{n}{100k_{i,j}}\rceil}$ and $\mathcal{L}_{\lceil\frac{n}{100k_{i,j}}\rceil}$ are (see Figure 6)
$a^{\prime}:$ $=\left(\frac{-\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan(\theta_{j})}-\frac{\ell^{1/96}n^{2/3}}{2\sqrt{2}},\frac{-\lceil\frac{n}{100k_{i,j}}\rceil\tan(\theta_{j})}{1+\tan(\theta_{j})}+\frac{\ell^{1/96}n^{2/3}}{2\sqrt{2}}\right),$
$b^{\prime}:$ $=\left(\frac{-\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan(\theta_{j})}+\frac{\ell^{1/96}n^{2/3}}{2\sqrt{2}},\frac{-\lceil\frac{n}{100k_{i,j}}\rceil\tan(\theta_{j})}{1+\tan(\theta_{j})}-\frac{\ell^{1/96}n^{2/3}}{2\sqrt{2}}\right),$
$c^{\prime}:$ $=\left(\frac{\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan(\theta_{j})}-\frac{\ell^{1/96}n^{2/3}}{2\sqrt{2}},\frac{\lceil\frac{n}{100k_{i,j}}\rceil\tan(\theta_{j})}{1+\tan(\theta_{j})}+\frac{\ell^{1/96}n^{2/3}}{2\sqrt{2}}\right),$
$d^{\prime}:$ $=\left(\frac{\lceil\frac{n}{100k_{i,j}}\rceil}{1+\tan(\theta_{j})}+\frac{\ell^{1/96}n^{2/3}}{2\sqrt{2}},\frac{\lceil\frac{n}{100k_{i,j}}\rceil\tan(\theta_{j})}{1+\tan(\theta_{j})}-\frac{\ell^{1/96}n^{2/3}}{2\sqrt{2}}\right).$
Then
$\psi(a)-\psi(a^{\prime})=-\frac{2\lceil\frac{n}{100k_{i,j}}\rceil\left(\tan(\theta_{i})-\tan(\theta_{j})\right)}{(1+\tan(\theta_{i}))(1+\tan(\theta_{j}))}+\frac{1}{\sqrt{2}}\left(\ell^{1/96}n^{2/3}-n^{2/3}\right).$
from the above discussion, this difference is much larger than $(\ell^{1/97}-1)n^{2/3}$. Similarly, we compute $\psi(b^{\prime})-\psi(b),\psi(c)-\psi(c^{\prime})$ and $\psi(d^{\prime})-\psi(d)$ to conclude the following:
If $\exists u\notin\mathsf{R}_{i}^{\ell}\text{ and }\exists\gamma_{i}\in A_{i}:|\phi(u)+n|\leq\lceil\frac{n}{100k_{i,j}}\rceil,\Gamma_{u}^{\gamma_{i}}$ intersects $W_{i,j}$ then, as $A_{i}$ is of length $\frac{1}{n^{1/3}}$, $\Gamma_{u}^{\gamma_{i}}$ will have transversal fluctuation at least $\sim(\ell^{1/97}-1)n^{2/3}$ from the time level $-n+\lceil\frac{n}{100k_{i,j}}\rceil$ to somewhere in between $-\lceil\frac{n}{100k_{i,j}}\rceil$ and $\lceil\frac{n}{100k_{i,j}}\rceil$ (this is precisely the reason for the choice of the direction of $W_{i,j}$ (see Figure 6). By an easy planarity argument now the probability of this event can be bounded by the desired upper bound as in the statement of the lemma. ∎
@@
For sufficiently large $n,\ell$ there exist constants $C,c>0$ (depending only on $\varepsilon$) such that
@@
This lemma is already proved in [BBB23, Lemma 3.1], the reader can refer to this proof. We briefly describe the idea here. We consider two points $a$ and $b$ with $\psi(a)<\psi(b)$ on $\mathcal{L}_{-n-\lceil\frac{n}{100k_{i,j}}\rceil}$ which are $2\ell^{1/96}n^{2/3}$ distance away (in the space direction) from the end points of $R_{i}^{\ell}$. Now, by transversal fluctuation estimate we can say that $\Gamma_{a}^{\beta_{i}}$ and $\Gamma_{b}^{\alpha_{i}}$ always stay $\ell^{1/96}n^{2/3}$ distance away (in the space direction) from the rectangle $R_{i}^{\ell}$ with probability at least $1-e^{-c\ell^{3/96}}$. On this event, every geodesics starting in $R_{i}^{\ell}$ will be sandwiched between $\Gamma_{a}^{\beta_{i}}$ and $\Gamma_{b}^{\alpha_{i}}$. So, on this event by planarity argument the event $\mathcal{G}_{i}$ will imply large transversal fluctuation of either $\Gamma_{a}^{\beta_{i}}$ or $\Gamma_{b}^{\alpha_{i}}$. This gives us the desired upper bound. ∎
@@
Note that on the second event in the right hand side there are more than $\ell^{1/3}$ distinct equivalence classes of geodesics starting from $\mathcal{J}_{i}^{\prime}$ and ending at $\mathcal{J}_{i}^{\prime\prime}$. Using Proposition 6.4, for $\ell<n^{0.03}$ sufficiently large (depending on $\varepsilon$) we have
@@
Same argument shows for $\ell<n^{0.03}$ sufficiently large (depending on $\varepsilon$) we have
@@
$\mathbb{P}\left(\max I_{u,v,\gamma_{i},\gamma_{j}}\geq\frac{\ell^{1/3}n\log|k_{i,j}|}{k_{i,j}^{3}}\right)\leq\mathbb{P}(\mathcal{F}_{i}\cup\mathcal{G}_{i}\cup\mathcal{F}_{j}\cup\mathcal{G}_{j})+$
($\mathbb{P}\left(\left\{\max I_{u,v,\gamma_{i},\gamma_{j}}\geq\frac{\ell^{1/3}n\log|k_{i,j}|}{k_{i,j}^{3}}\right\}\cap(\mathcal{F}_{i}\cup\mathcal{G}_{i}\cup\mathcal{F}_{j}\cup\mathcal{G}_{j})^{c}\right).$
@@
On the event $\left\{\max I_{u,v,\gamma_{i},\gamma_{j}}\geq\frac{\ell^{1/3}n\log|k_{i,j}|}{k_{i,j}^{3}}\right\}\cap(\mathcal{F}_{i}\cup\mathcal{G}_{i}\cup\mathcal{F}_{j}\cup\mathcal{G}_{j})^{c}$, we want to estimate the maximum size of the intersection of geodesics inside $W_{i,j}$ that are starting from $\mathcal{J}_{i}^{\prime}$ (resp. $\mathcal{J}_{j}^{\prime}$) and ending at $\mathcal{J}_{i}^{\prime\prime}$ (resp. $\mathcal{J}_{j}^{\prime\prime}$). For this we use Lemma 6.6. Hence, we have that there exist constants $C,c$ (depending on $\varepsilon$) such that for large enough $\ell$ and $n$
($\mathbb{P}\left(\left\{\max I_{u,v,\gamma_{i},\gamma_{j}}\geq\frac{\ell^{1/3}n\log|k_{i,j}|}{k_{i,j}^{3}}\right\}\cap(\mathcal{F}_{i}\cup\mathcal{G}_{i}\cup\mathcal{F}_{j}\cup\mathcal{G}_{j})^{c}\right)\leq Ce^{-c\ell^{1/3}}.$
Hence, combining the above we get that there exists a constant $C>0$ (depending on $\varepsilon$) such that for all large enough $n$ we have
@@
hence, from (3.7) we have that there exists $C_{2}>0$ such that for sufficiently large $n$
@@
We now go back to (3.4). From (3.6) and (3.13) we have that there exists constants $c_{1},C_{2}>0$, large $M$ such that for sufficiently large $n$(all depending only on $\varepsilon$)
@@
As $C,c$ depend only on $\varepsilon$ (as we have chosen all the interval length to be $\frac{1}{n^{1/3}}$) we can choose $M$ large enough so that there exists $c>0$ such that for sufficiently large $n$